Systems Staff Meeting Notes 01/07/02

Systems Staff Meeting Notes
January 7, 2002

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Suzy Covey, Winston Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart, Ying Tang
  1. Hula girl, tiki torches in Library East – it’s a performance art installation by an artist from St. Augustine. There will also be some ancient calligraphy on the walls and a likeness of Samuel Morris set up in the front broom closet. Speakers will be set up – expect noise – showing to last the duration of two months.
  2. Vault inspection – PPD – Bill’s been dosed twice, wants no more. He should be notified if onsite inspection is imminent.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Orders – will order machines again this month – 20 to 40 at a time.
  2. Will also order a SAN backup server for use as the backup of SmathersNT2R.
  3. Education – check on the electrical power status – replacement of power extension. They want better support – they’re adding a cappuccino machine.
Support Issues and Liaisons
  1. W2K rollout – has rolled in West, progress in AFA. MSL scheduled for tomorrow at 8:30AM. Security settings review – they have 2 images – E3000 & E4200. We have our own GX110 image that they want to modify & make their own security. This is okay. The classroom in West is in a tizzy – it is not upgraded to W2K yet. They are in the process of sorting out what the principle problem is – people will have to have an ID – we’ve provided an automatic way to do it – people will be responsible for the network security. W2K not upgraded on staff machines in LAC, Education, & Judiaca. Personnel is not backed up – the liaison needs instruction. Will showed Cheryl how & why, but they are not backed up yet. Resource Services has volunteered to help in LAC & Education.
  2. Enforcing contact channels – liaisons going through SYSHELP. Bill has informed the Directors this morning about the need for departments to go through the proper channels.
  1. Current developments – new BadTrans variant – W32 BadTrans-mm.enc. JPG not detected as a virus. The Outlook mailbox was wiped on the server side after sending it to all in the address book. In theory, this will not bother us – we block JPEG. Variants of Majister are still hitting. Going to an outside server can bring it back in. Cynthia got a notification from our virus server scan while out at a wallpaper site: Johnny X. TrendScan is located in Japan – they should get the cure out. Ancient version of virus checker should be a flag problem. Personal firewall – Zone Alarm – can cut virus checker updates off.
  2. Facilities Security System – Bill received a message from Rob Roberts to hook up to the Campus Police. VPN & equipment on hand – need to be careful granting access to outside groups – their access should be controlled. Campus police did not want to be a part of this.
  3. Mailbox backups – lack of…trying again. Have placed a call to Veritas Software to obtain a fix. Has been making attachments fail in email.
  4. W2K server – next week. Central defrag – soon.
  5. Email for Public areas – Geoffrey excluded it from the Proxy – West machines are on the death list – swapping for kiosk machines. Will not bring up W2K on these. Revisiting the ideal of doing email – tool worthy or not? Should be on all if needed or banned on all if not. For the time being, they will be excluded from the proxy. Almost every email has Web interface. Report of HotMail access even going through the proxy.
Software Projects Status
  1. ILLIAD – problem with patron registers. Enters email address. If outside of, it doesn’t work. May be a SMTP problem – open mail relay. Internet – not to permit mail servers as relays – ran into this problem with Eshare for RefExpress – may have to permit NT15 to be on the list for open relay. server doesn’t require authentication at the moment. All automatic mailings will have to deal with this – they must be authenticated.
  2. FilmLog – ongoing. Cynthia finished interim updates for patron registrations. Winston discussed this with John Ingram – he’s very interested in getting it going. Patron and staff side – making prototype. Back up verification for Patron ID.
  3. Special Collections – working on Web migration – rebuilt folder structure – will have to rebuild groups – trying to anticipate AD, maybe test next week – content site.
DP Coordinator
  1. Debra is out today. Students should be starting to work – Kristen called to say she will not be in tomorrow, but will start on Wednesday.

Back to the Systems Homepage

Last updated January 14, 2002
by Debra Harris